### v1

# Interpolated time-Hölder regularity of solutions of fully nonlinear parabolic equations

Abstract.
We show by the maximum principle that a quantitative estimate of the Hölder continuity of a solution of a fully nonlinear parabolic equation with respect to the spatial variable $x\in\mathbb{R}^{n}$, with exponent $\alpha\in(0,1]$, implies an estimate of its Hölder continuity in the time variable with the (natural) exponent $\alpha/2$. This holds without assuming any Hölder continuity of the data of the equation. We apply this nonvariational approach to prove interior Schauder estimates for a special class of fully nonlinear parabolic Isaacs equations, providing an Evans-Krylov result for the model equation $\min\{\inf_{\beta}L_{\beta}u,\sup_{\gamma}L_{\gamma}u\}-\partial_{t}u=0$, where $L_{\beta},L_{\gamma}$ are linear operators with possibly variable Hölder coefficients. Along the way, we provide a short survey on the regularity theory for fully nonlinear parabolic equations of second order.
Key words and phrases:
Evans-Krylov theorem, Fully nonlinear parabolic equations, Isaacs equations
2020 Mathematics Subject Classification:
35B65
The author is member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). He was partially supported by the INdAM-GNAMPA project 2024, the King Abdullah University of Science and Technology (KAUST) project CRG2021-4674 “Mean-Field Games: models, theory and computational aspects” and by the project funded by the EuropeanUnion – NextGenerationEU under the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.1 - Call PRIN 2022 No. 104 of February 2, 2022 of Italian Ministry of University and Research; Project 2022W58BJ5 (subject area: PE - Physical Sciences and Engineering) “PDEs and optimal control methods in mean field games, population dynamics and multi-agent models”
1. Introduction
In this note we provide a maximum principle approach to show that the knowledge of a Hölder bound in space with exponent $\alpha\in(0,1]$ of solutions to fully nonlinear parabolic equations with bounded coefficients of the form
$F(x,t,Du,D^{2}u)-\partial_{t}u=f(x,t)$
implies an a priori bound on its (natural) time-Hölder norm of order $\alpha/2$. The proof builds upon the construction of a suitable barrier function and the comparison principle for fully nonlinear parabolic equations. This method was sometimes used in the literature of linear PDEs by S. Kruzhkov [20, 21], later refined in [16] and in [2] for the one-dimensional Porous Medium flow. Remarkably, no Hölder regularity in time of the data of the equation are required to perform this step.
We apply this procedure to study interior maximal $C^{\alpha}$ smoothness (i.e. local $C^{2,\alpha}$ regularity with respect to the parabolic distance) of solutions to the following fully nonlinear parabolic model
(1) $F(D^{2}u)-\partial_{t}u=0\text{ in }Q_{1}:=B_{1}\times(-1,0]\subseteq\mathbb{R}^{n+1},$
where we assume
(2) $\begin{cases}F(M)=\min\{F^{\cap}(M),F^{\cup}(M)\},\ \forall M\in\mathcal{S}_{n},\\ F(0)=0,\ F^{\cap},F^{\cup}\text{ are uniformly elliptic},\\ F^{\cap}\text{ is concave in $M$},\ F^{\cup}\text{ is convex in $M$}.\end{cases}$
The uniform ellipticity for an operator $G=G(M):\mathrm{Sym}_{n}\to\mathbb{R}$, $\mathrm{Sym}_{n}$ being the space of $n\times n$ symmetric matrices, reads as
$\lambda\|N\|\leq G(M+N)-G(M)\leq\Lambda\|N\|,\ \forall M\in\mathrm{Sym}_{n},\ N\geq 0,\ \lambda\leq\Lambda.$
Note that the uniform ellipticity of $F^{\cap},F^{\cup}$ implies that $F$ itself is uniformly elliptic. In this case we will call (1) uniformly parabolic. The result applies for instance to the evolution PDE
(3) $\min\left\{\inf_{a\in\mathcal{A}}L_{a}u,\sup_{b\in\mathcal{B}}L_{b}u\right\}-\partial_{t}u=0,$
where $\mathcal{A},\mathcal{B}$ are arbitrary sets and $L_{a},L_{b}$ are linear operators of the form $L_{k}u=a_{ij}^{k}\partial_{ij}u+c_{k}$, $c_{k}=L_{k}0\in\mathbb{R}$, $a_{ij}^{k}$ with constant coefficients. This class can be written in the Isaacs form $\inf_{\sigma\in\Sigma}\sup_{\theta\in\Theta}L_{\sigma\theta}u-\partial_{t}u=0$, where the convex operators $\left(\sup_{\theta\in\Theta}L_{\sigma\theta}u\right)_{\sigma\in\Sigma}$ are all linear except for at most one $\sigma$. A prototype equation is the flow driven by the 3-operator
$F_{3}(D^{2}u)-\partial_{t}u=\min\{L_{1}u,\max\{L_{2}u,L_{3}u\}\}-\partial_{t}u,$
where $L_{1},L_{2},L_{3}$ are linear uniformly elliptic operators with constant coefficients. This provides a time-dependent counterpart of a result by L. Caffarelli and X. Cabré [5] and it can be regarded as an Evans-Krylov type result for a (time-dependent) nonconvex/nonconcave equation. We recall that by the Evans-Krylov property we mean the boost of regularity
$C^{2,1}\rightarrow C^{2+\alpha,1+\alpha/2},$
see Theorem 6.1 of [8] for the elliptic result.
The outline of the proof of the maximal parabolic $C^{\alpha}$ regularity under the assumption (2) is the following: one first proves that for the homogeneous equation $F(D^{2}u)-\partial_{t}u=0$ we have
$\partial_{t}u,Du\in C^{\alpha,\alpha/2}\text{ (in particular $Du\in C_{t}^{\alpha/2}$)}$
for some small universal $\alpha\in(0,1)$, by the Krylov-Safonov theory [25] (here $C_{t}^{\alpha/2}$ denotes the Hölder continuity in time, uniformly in space). Indeed, the first derivative (or its incremental quotient), both in time and space, satisfies a linear nondivergence form equation, since the equation is invariant in space-time. This step does not require any convexity or concavity assumptions on $F$, and it holds for continuous viscosity solutions of any uniformly parabolic equation, see Theorem 4.4. Then, for each time slice, which we freeze and do not display, we write
$F(D^{2}u)=g$
where $g=\partial_{t}u\in C^{\alpha,\alpha/2}$. This implies by the stationary (space) regularity of [5]
$u\in C^{2,\alpha}_{x}$
at each time layer. To reach parabolic $C^{2+\alpha,1+\alpha/2}$ regularity of the solution to the parabolic equation (and in particular $\alpha/2$ time-regularity of second derivatives) we use the previous idea based on the maximum principle method at a higher order level. More precisely, exploiting the information $D^{2}u\in C^{\alpha}_{x}$, we improve the time-regularity of the gradient showing that
$Du\in C_{t}^{\frac{1+\alpha}{2}}$
via the comparison principle. We then exploit an interpolation argument between $Du\in C_{t}^{(1+\alpha)/2}$ and $D^{2}u\in C^{\alpha}_{x}$ to conclude that $D^{2}u\in C^{\alpha/2}_{t}$, proving thus the claimed $C^{2+\alpha,1+\alpha/2}$ regularity. Note that a direct interpolation with the sole information $Du\in C^{\alpha/2}_{t}$ provides a worse Hölder regularity exponent than the expected regularity in $C^{2+\alpha,1+\alpha/2}$, which is however enough in general to reach classical smoothness.
In the case of the heat equation, a maximum principle approach was used to prove the time-Hölder regularity of the Hessian in [19] via an argument of A. Brandt [3].
Notably, our method provides a new proof of the parabolic $C^{2,\alpha}$ regularity for fully nonlinear, concave, uniformly parabolic equations using the stationary result and the maximum principle, and can be generalized to any parabolic equation that admits $C^{2,\alpha}$ interior elliptic estimates, see Section 4. A comparison with other approaches to study the parabolic regularity of fully nonlinear equations is in Remark 4.7.
The study of these estimates are motivated by the challenging open question of finding some special structures of $F$ guaranteeing classical smoothness of solutions when the parabolic dimension $n+1\geq 4$, cf. the introduction of [17] and Section 3 for references on the subject.
Following [5], we mention that these interior estimates could be used to prove an existence result for the Dirichlet problem by a refinement of the method of continuity based on weighted interior bounds. This approach was first introduced by J. H. Michael [30] for linear stationary PDEs and extended to linear parabolic equations in [34]. We do not pursue this direction here, but just recall that the application of the classical method of continuity requires estimates up to the boundary along with a smooth operator, namely $F$ must be at least of class $C^{1}$. Moreover, we emphasize that the process of smoothing fully nonlinear operators is known only for concave/convex operators, and hence one needs a different approach for other classes of non-smooth operators.
Plan of the paper. Section 2 provides a general result at the zero-th order level saying that one can use the equation and the space Hölder regularity to deduce the time-Hölder regularity via the maximum principle. Section 3 provides a brief survey on the known literature of the $C^{2,\alpha}$ solvability of fully nonlinear parabolic equations. Section 4 is devoted to the application of the maximum principle approach to obtain $C^{2+\alpha,1+\alpha/2}$ regularity for a time-dependent counterpart of an Isaacs model introduced by L. Caffarelli and X. Cabré.
Notation
We denote by $B_{r}(x)$ the ball of center $x$ and radius $r$. When $x_{0}=0$ we simply write $B_{r}$.
$Q_{r}(x,t)$ is the parabolic cylinder $B_{r}(x)\times(t-r^{2},t)$, and we write $Q_{r}$ when $(x,t)=(0,0)$.
We denote, given $\Omega$ an open set, by $\partial_{\mathrm{par}}(\Omega\times(a,b))$ the parabolic boundary of a set.
We denote by $|u|_{0;\Omega}$ the sup-norm of $u$, i.e. $\|u\|_{C(\Omega)}$ (both in the elliptic and the parabolic case).
Let $Q\subset\Omega\times(0,T)$ and $\alpha\in(0,1)$. We denote by $d((x,t),(y,s))=|x-y|+|t-s|^{\frac{1}{2}}$ the standard parabolic distance and
  • $C^{\alpha,\alpha/2}(Q)$, $\alpha\in(0,1]$, the space of those $u:Q\to\mathbb{R}$ such that
    $\|u\|_{C^{\alpha,\alpha/2}(Q)}:=|u|_{0;Q}+[u]_{C^{\alpha,\alpha/2}(Q)}=\|u\|_{L^{\infty}(Q)}+\sup_{\begin{subarray}{c}(x,t),(y,s)\in Q,\\ (x,t)\neq(y,s)\end{subarray}}\frac{|u(x,t)-u(y,s)|}{d^{\alpha}((x,t),(y,s))};$
  • $C^{1+\alpha,(1+\alpha)/2}(Q)$ the space of functions $u$ whose spatial gradient exists classically and equipped with the norm
    $\|u\|_{C^{1+\alpha,(1+\alpha)/2}(Q)}:=|u|_{0;Q}+|Du|_{0;Q}+\sup_{\begin{subarray}{c}(x,t),(y,s)\in Q,\\ (x,t)\neq(y,s)\end{subarray}}\frac{|u(x,t)-u(y,s)|}{d^{1+\alpha}((x,t),(y,s))}.$
    In particular, any $u\in C^{1+\alpha,(1+\alpha)/2}(Q)$ is such that each component of $Du$ belongs to $C^{\alpha,\alpha/2}(Q)$ and $u$ is Hölder continuous with exponent $(1+\alpha)/2$ in the time variable;
  • $C^{2+\alpha,1+\alpha/2}(Q)$ the space of functions $u$ such that
    $\|u\|_{C^{2+\alpha,1+\alpha/2}(Q)}:=|u|_{0;Q}+\sum_{i=1}^{n}\|u_{x_{i}}\|_{C^{1+\alpha,(1+\alpha)/2}(Q)}+\|\partial_{t}u\|_{C^{\alpha,\alpha/2}(Q)}$
    This is equivalent to say that $D^{2}u$ belongs to $C^{\alpha,\alpha/2}(Q)$ and $\partial_{t}u$ belongs to $C^{\alpha,\alpha/2}(Q)$: it is a consequence of Remark 8.8.7 in [23]. In this case we can consider the space $C^{2+\alpha,1+\alpha/2}(Q)$ equipped with the norm
    $\|u\|_{C^{2+\alpha,1+\alpha/2}(Q)}:=|u|_{0;Q}+|Du|_{0;Q}+\|D^{2}u\|_{C^{\alpha,\alpha/2}(Q)}+\|\partial_{t}u\|_{C^{\alpha,\alpha/2}(Q)}.$
    For more properties on these spaces we refer to [23]. Furthermore, we denote by
    $\|u\|_{C^{2,1}(Q)}:=\sum_{2i+j\leq 2}|\partial_{t}^{i}D_{x}^{j}u|_{0;Q}.$
    We will also use the equivalence (see e.g. p. 120 of [23]) between the Hölder seminorm $[u]_{C^{\alpha,\alpha/2}(Q_{1})}$ previously defined and the seminorm
    $[u]^{\prime}_{C^{\alpha,\beta}(Q_{1})}=\sup_{\begin{subarray}{c}x,y\in B_{1},\\ x\neq y\end{subarray}}\frac{|u(x,t)-u(y,t)|}{|x-y|^{\alpha}}+\sup_{\begin{subarray}{c}t,s\in(-1,0),\\ t\neq s\end{subarray}}\frac{|u(x,t)-u(x,s)|}{|t-s|^{\frac{\alpha}{2}}}.$
2. Time-Hölder regularity through the maximum principle
We consider the operator
(4) $F(x,t,Du,D^{2}u)-\partial_{t}u=f(x,t)\text{ in }Q=\Omega\times(0,T],$
where $\Omega$ is an open connected subset of $\mathbb{R}^{n}$, and $T$ is some fixed positive constant. We assume that for $X,Y\in\mathrm{Sym}_{n}$, $p,q\in\mathbb{R}^{n}$ and $(x,t)\in Q$
(5) $\mathcal{M}^{-}_{\lambda,\Lambda}(M-N)-K|p-q|\leq F(x,t,p,M)-F(x,t,q,N)\leq\mathcal{M}^{+}_{\lambda,\Lambda}(M-N)+K|p-q|$
with $F(x,t,0,0)=0$ and $\mathcal{M}^{\pm}_{\lambda,\Lambda}$ are the Pucci’s extremal operators with constants $\lambda\leq\Lambda$ defined by
$\mathcal{M}^{+}_{\lambda,\Lambda}(M)=\sup_{\lambda I_{n}\leq A\leq\Lambda I_{n}}\mathrm{Tr}(AM)=\Lambda\sum_{e_{k}>0}e_{k}+\lambda\sum_{e_{k}<0}e_{k},$
$\mathcal{M}^{-}_{\lambda,\Lambda}(M)=\sup_{\lambda I_{n}\leq A\leq\Lambda I_{n}}\mathrm{Tr}(AM)=\lambda\sum_{e_{k}>0}e_{k}+\Lambda\sum_{e_{k}<0}e_{k},$
where $e_{k}=e_{k}(M)$ are the eigenvalues of $M$.
We denote by $\underline{\mathcal{S}}(\lambda,\Lambda,f)$ the space of continuous functions $u$ in $\Omega$ that are subsolutions of the evolutive Pucci’s maximal equation
$\mathcal{M}^{+}_{\lambda,\Lambda}(D^{2}u)-\partial_{t}u=f(x,t)\text{ in the viscosity sense in $\Omega$},$
Similarly, $\overline{\mathcal{S}}(\lambda,\Lambda,f)$ will denote the space of continuous functions $u$ in $\Omega$ that are supersolutions to the Pucci’s minimal equation
$\mathcal{M}^{-}_{\lambda,\Lambda}(D^{2}u)-\partial_{t}u=f(x,t)\text{ in the viscosity sense in $\Omega$},$
Solutions of fully nonlinear uniformly elliptic equations $F(D^{2}u)=0$ belong to the class $\mathcal{S}(\lambda,\Lambda,0)=\overline{\mathcal{S}}(\lambda,\Lambda,0)\cap\underline{\mathcal{S}}(\lambda,\Lambda,0)$.
We further assume that the right-hand side $f\in C(Q)$ satisfies
$|f(x,t)|\leq C_{f}$
for some positive constant $C_{f}$. The result below applies to solutions of the model PDE in nondivergence form
$F(x,t,D^{2}u)+b(x,t)\cdot Du-\partial_{t}u=f(x,t)$
as well as to the viscous Hamilton-Jacobi equation
$\mathrm{Tr}(A(x,t)D^{2}u)+b(x,t)|Du|-\partial_{t}u=f(x,t)$
in the case of continuous and bounded coefficients $A,b,f$. Let $\Omega^{\prime}$ be an open connected subset of $\Omega$ such that for some $d>0$
$d\leq\mathrm{dist}(\Omega^{\prime},\partial\Omega)$
and set $Q^{\prime}=\Omega^{\prime}\times(0,T]$. The next proof shows by the maximum principle the (quantitative) Hölder regularity in time of order $\alpha/2$ by means of the comparison principle and the construction of a suitable barrier, knowing a priori a quantitative Hölder bound in the space variable. It can be regarded as an a priori estimate.
Theorem 2.1.
Let $u$ be a $C^{2,1}$ solution to (4) in $Q$ such that for all $(x,t),(y,t)\in Q$ we have the estimate
(6) $|u(x,t)-u(y,t)|\leq C_{1}|x-y|^{\alpha}$
for some $\alpha\in(0,1]$ and some constant $C_{1}>0$. Then there exists a positive constant $\delta$ depending on $C_{f}$ and $d$, where $d\leq\mathrm{dist}(\Omega^{\prime},\partial\Omega)$, and a positive constant $C_{2}$, depending on $C_{f},C_{1},K,n,\Lambda$ and $d$, such that
$|u(x,t)-u(x,t_{0})|\leq C_{2}|t-t_{0}|^{\frac{\alpha}{2}}$
for all $(x,t),(x,t_{0})\in Q^{{}^{\prime}}$ with $|t-t_{0}|<\delta$.
Proof.
Let $(x_{0},t_{1}),(x_{0},t_{2})$ be arbitrary points in $Q^{\prime}$. We denote for $t_{1}<t_{2}\leq T$ and $R\in(0,d)$ by
$\mathbb{Q}^{t_{1},t_{2}}_{x_{0},R}:=\{(x,t)\in\mathbb{R}^{n+1}:|x-x_{0}|<R,t_{1}<t\leq t_{2}\}\subseteq Q$
a cylinder in $\mathbb{R}^{n+1}$. We consider in $\mathbb{Q}^{t_{1},t_{2}}_{x_{0},R}$ the functions
$\varphi^{\pm}(x,t)=[C_{f}+2\Theta R^{-2}(n\Lambda+KR)](t-t_{1})+\Theta R^{-2}|x-x_{0}|^{2}+C_{1}R^{\alpha}\pm[u(x,t)-u(x_{0},t_{1})],$
where
$\Theta=\sup_{t\in[t_{1},t_{2}]}|u(x_{0},t)-u(x_{0},t_{1})|.$
We have by (5)
$F[\varphi^{\pm}]$ $:=F(x,t,D\varphi^{\pm},D^{2}\varphi^{\pm})-\partial_{t}\varphi^{\pm}=-C_{f}-2\Theta R^{-2}(n\Lambda+KR)+F(x,t,D\varphi^{\pm},D^{2}\varphi^{\pm})$
$\leq-C_{f}-2\Theta R^{-2}(n\Lambda+KR)+F(x,t,0,0)+\mathcal{M}^{+}_{\lambda,\Lambda}(D^{2}\varphi^{\pm})+K|D\varphi^{\pm}|$
$\leq-C_{f}-2\Theta R^{-2}(n\Lambda+KR)+2\Theta R^{-2}n\Lambda+2\Theta R^{-2}K|x-x_{0}|\pm f(x,t)\leq 0.$
Note that on the layer $t=t_{1}$ and $|x-x_{0}|\leq R$ we have by (6)
$\varphi^{\pm}(x,t)=\Theta R^{-2}|x-x_{0}|^{2}+C_{1}R^{\alpha}+[u(x,t_{1})-u(x_{0},t_{1})]\geq\Theta R^{-2}|x-x_{0}|^{2}+C_{1}R^{\alpha}-C_{1}R^{\alpha}\geq 0.$
On the lateral surface of the cylinder $|x-x_{0}|=R$ and $t\in(t_{1},t_{2}]$ we have by (6) and the definition of $\Theta$
$\varphi^{\pm}(x,t)$ $=[C_{f}+2\Theta R^{-2}(n\Lambda+KR)](t-t_{1})+\Theta+C_{1}R^{\alpha}\pm[u(x,t)-u(x_{0},t_{0})]$
$\geq C_{1}R^{\alpha}\pm[u(x,t)-u(x_{0},t)]+\Theta\pm[u(x_{0},t)-u(x_{0},t_{0})]\geq 0.$
Thus $F[\varphi^{\pm}]\leq 0$ in $\mathbb{Q}^{t_{1},t_{2}}_{x_{0},R}$ and $\varphi^{\pm}\geq 0$ on $\partial_{\mathrm{par}}\mathbb{Q}^{t_{1},t_{2}}_{x_{0},R}$. By the comparison principle [10, Theorem 8.2] we have $\varphi^{\pm}\geq 0$ in $\mathbb{Q}^{t_{1},t_{2}}_{x_{0},R}$. More precisely
$\mp[u(x_{0},t)-u(x_{0},t_{0})]\leq C_{1}R^{\alpha}+[C_{f}+2\Theta R^{-2}(n\Lambda+KR)](t_{2}-t_{1})$
for any $t\in(t_{1},t_{2}]$. Hence, passing to the supremum on the left-hand side
(7) $\Theta\leq C_{1}R^{\alpha}+C_{f}(t_{2}-t_{1})+\frac{1}{2}\Theta[4R^{-2}(n\Lambda+KR)](t_{2}-t_{1}).$
We denote by $R^{*}$ the positive root of the quadratic equation
$4R^{-2}(n\Lambda+KR)(t_{2}-t_{1})=1\implies R^{2}-4KR(t_{2}-t_{1})-4n\Lambda(t_{2}-t_{1})=0,$
namely
$R^{*}=2K(t_{2}-t_{1})+2\sqrt{K^{2}(t_{2}-t_{1})^{2}+n\Lambda(t_{2}-t_{1})}.$
Note that if $t_{2}<t_{1}+\delta$,
$\delta=\frac{d^{2}}{4(n\Lambda+dK)},$
we have $R^{*}<d$, so $R^{*}$ is an admissible value. Replacing $R$ with $R^{*}$ in (7) we conclude
$\Theta\leq C_{1}(R^{*})^{\alpha}+C_{f}(t_{2}-t_{1})+\frac{1}{2}\Theta.$
This implies
$\frac{1}{2}\Theta$ $\leq C_{1}\left\{2K(t_{2}-t_{1})+2\sqrt{K^{2}(t_{2}-t_{1})^{2}+n\Lambda(t_{2}-t_{1})}\right\}^{\alpha}+C_{f}\delta^{1-\frac{\alpha}{2}}(t_{2}-t_{1})^{\frac{\alpha}{2}}$
$\leq C_{1}\left\{(2K)^{\alpha}(t_{2}-t_{1})^{\alpha}+2^{\alpha}[K^{2}(t_{2}-t_{1})^{2}+n\Lambda(t_{2}-t_{1})]^{\frac{\alpha}{2}}\right\}+C_{f}\delta^{1-\frac{\alpha}{2}}(t_{2}-t_{1})^{\frac{\alpha}{2}}$
$\leq C_{1}\left\{(2K)^{\alpha}(t_{2}-t_{1})^{\alpha}+(2K)^{\alpha}(t_{2}-t_{1})^{\alpha}+(n\Lambda)^{\frac{\alpha}{2}}(t_{2}-t_{1})^{\frac{\alpha}{2}}\right\}+C_{f}\delta^{1-\frac{\alpha}{2}}(t_{2}-t_{1})^{\frac{\alpha}{2}}$
$\leq C_{1}(t_{2}-t_{1})^{\frac{\alpha}{2}}\left\{2(2K)^{\alpha}(t_{2}-t_{1})^{\frac{\alpha}{2}}+(n\Lambda)^{\frac{\alpha}{2}}\right\}+C_{f}\delta^{1-\frac{\alpha}{2}}(t_{2}-t_{1})^{\frac{\alpha}{2}}$
$\leq C_{1}(t_{2}-t_{1})^{\frac{\alpha}{2}}\left\{2(2K)^{\alpha}K^{-1}d^{\frac{\alpha}{2}}+(n\Lambda)^{\frac{\alpha}{2}}\right\}+C_{f}\delta^{1-\frac{\alpha}{2}}(t_{2}-t_{1})^{\frac{\alpha}{2}}.$
We thus end up with the inequality
$\Theta\leq C_{2}(t_{2}-t_{1})^{\frac{\alpha}{2}}$
for a constant $C_{2}$ depending on $n,C_{f},C_{1},K,\Lambda,d$. ∎
Remark 2.2.
It is worth noting that we only need the one-side comparison with the Pucci’s maximal operator $\mathcal{M}^{+}$ to conclude the argument. We can thus weaken the uniform ellipticity (5) imposing only the second inequality along with suitable ellipticity assumptions guaranteeing the validity of the comparison principle.
3. A quick survey on the (higher) regularity theory for fully nonlinear parabolic equations
In this section we survey on regularity results at the level of Hölder spaces for fully nonlinear parabolic equations, focusing on those of the form
(8) $F(D^{2}u)-\partial_{t}u=0\text{ in }B_{1}\times(-1,0]\subset\mathbb{R}^{n+1}.$
We will in particular concentrate on structural conditions, possibly involving the dimension, guaranteeing classical solutions. The interested reader can find a complete updated account in the book [15] for the elliptic theory. We do not aim at discussing here $W^{2,1}_{q}$ regularity of solutions and we refer to [24] for more details and to the next section for the main results on the low-regularity theory.
The simplest result for (8) says that solutions to uniformly parabolic equations in 3D (here we denote by $D$ the dimension in space-time) are always classical without any other assumption on $F$, see [1, 17, 33] and the references therein. Its elliptic analogue was proved by L. Nirenberg (in the elliptic dimension $n=2$) [15, Theorem 4.9], and its proof, based on the De Giorgi-Nash-Moser theory applied to the equation solved by second derivatives, does not seem to generalize to parabolic equations. It is also important to mention a result of S. Kruzhkov showing that the same result holds in the 2D case, cf. [20] and Section XIV.7 in [27].
Theorem 3.1.
Assume that $F$ is uniformly elliptic and $n=2$. Then, viscosity solutions to (8) (in dimension $n+1=3$) are always classical and belong to $C^{2+\alpha,1+\alpha/2}_{\mathrm{loc}}$ for some small universal $\alpha\in(0,1)$. It holds
$\|u\|_{C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})}.$
A second result holds under Cordes-type assumptions on the ellipticity in any space dimension, see Theorem 6.4 in [17] or [11].
Theorem 3.2.
Assume that $F$ is uniformly elliptic and $\frac{\Lambda}{\lambda}\leq 1+\delta$ for a small (explicit) constant $\delta$ depending only on $n$. Then viscosity solutions to (8) are always classical (in any dimension) and belong to $C^{2+\alpha,1+\alpha/2}_{\mathrm{loc}}$ for some small universal $\alpha$. In addition, we have the regularity estimate
$\|u\|_{C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})}.$
We also mention that $C^{2+\alpha,1+\alpha/2}$ estimates can be obtained without any concavity assumptions for short time horizons, see e.g. Chapter 8 of [29], [28] and the references therein. This corresponds to an elliptic result proved in [14] and saying that $C^{2,\alpha}$ and higher order estimates for the equation $F(x,u,Du,D^{2}u)+\theta u=0$ can be achieved if the zero-th order coefficient $\theta$ is large.
A counterpart of the result by Evans-Krylov [13, 24] provides parabolic $C^{2,\alpha}$ estimates under concavity assumptions on $F$, cf. [36], see also [27, 22, 24] for smooth operators.
Theorem 3.3.
Assume that $F$ is uniformly elliptic and concave. Then viscosity solutions to (8) are always classical and belong to $C^{2+\alpha,1+\alpha/2}_{\mathrm{loc}}$ for some small universal $\alpha$. It holds
$\|u\|_{C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})}.$
Concavity on the second derivatives was recently weakened to the requirement of convexity of the superlevel sets, cf. [17]. The latter also contains some $C^{1,1}$ estimates under concavity-type conditions at infinity implying, among others, $W^{2,1}_{q}$ estimates for non-homogeneous equations. A recent account on the treatment of fully nonlinear parabolic equations with relaxed concavity conditions can be found in [24].
We should mention that slightly before the Evans-Krylov theory, the paper [32] provided an obstacle problem approach to reach parabolic $C^{2,\beta}$, $\beta$ small and universal, regularity for the model equation
$\max\{\partial_{t}u-L_{1}u,\partial_{t}u-L_{2}u\}=0$
(here $L_{1},L_{2}$ are uniformly elliptic linear operators) by reducing the problem to a variational inequality, an idea introduced by Brézis-Evans [4]. More references on the early regularity theory of fully nonlinear elliptic and parabolic equations can be found in [27, p. 382-384].
More recently, maximal $C^{\alpha}$ estimates were also achieved for flat solutions (i.e. with small $L^{\infty}$ norm) in [37] without concavity assumptions on $F$, extending a result due to O. Savin.
We also emphasize that counterexamples to the smoothness of solutions when $F$ is only uniformly parabolic can be found in the recent paper [33], which provides a time-dependent counterpart of the analysis in [31].
Different kind of results were studied in the paper [12], where the author provided a partial regularity result, again without imposing any concavity condition on $D^{2}u$, as initiated by Armstrong-Silvestre-Smart for the stationary problem, under the requirement that $F\in C^{1}$. It gives that a viscosity solution of a uniformly parabolic equation is $C^{2+\alpha,1+\alpha/2}$ on the complement of a closed set having Hausdorff dimension at most $n+1-\varepsilon$, where $\varepsilon$ depends on $\lambda,\Lambda,n$.
Finally, [9] studied the problem of the optimal parabolic regularity for the simple Bellman equation
$\max\{L_{1}u,L_{2}u\}-\partial_{t}u=0,$
providing a counterexample to the $C^{2,1}$ regularity of its solution. It seems a difficult open question to find the optimal regularity of solutions to general Bellman equations, even with constant coefficients, assuming only $F$ continuous and the uniform ellipticity. In this framework the best known elliptic and parabolic regularity in general dimension $n\geq 2$ is the Evans-Krylov $C^{2,\beta}$ regularity with $\beta$ small. This problem was raised in [4, Remark 4.2], see also [15], and the paper [6] highlighted that for the elliptic PDE $\max\{L_{1}u,L_{2}u\}=0$ in $\mathbb{R}^{2}$, viscosity solutions are of class $C^{2,1}$, while it remains open the general case $n\geq 3$.
Summarizing, for a uniformly parabolic equation of the form (8) with $F$ smooth we have the following results:
  • If $u\in C^{2,1}_{x,t}$, then $u\in C^{\infty}$ by the linear theory, see [17, Proposition 5.3];
  • $Du,\partial_{t}u\in C^{\alpha,\alpha/2}_{\mathrm{loc}}$ always by the Krylov-Safonov theory, see Theorem 4.4;
  • Solutions are always smooth for short times without concavity assumptions, see [28] and Chapter 8 of [29], or for arbitrary times when $F$ is a nonlinear function of the Laplacian, i.e. $F=f(\Delta u)$, see [35] and Section 8.5.2 of [29];
  • In the parabolic dimensions $D=2,3$, solutions are always smooth, see Theorem 3.1 and [20];
  • In any dimension, if $\frac{\Lambda}{\lambda}\leq 1+\delta(n)$, then the solutions are always smooth, see Theorem 3.2;
  • If $F$ is concave, then $u\in C^{\infty}$, see Theorem 3.3 and Corollary 8.7 in [17];
  • $u\in C^{2+\alpha,1+\alpha/2}_{\mathrm{loc}}$ in the case of flat solutions, see [37];
  • The assumption of concavity of $F$ can be weakened to convexity of superlevel sets, see [17];
  • Solutions of time-dependent Bellman equations cannot be $C^{2,1}$, see [9].
4. $C^{2+\alpha,1+\alpha/2}$ regularity for fully nonlinear parabolic Isaacs equations
Our main result is an Evans-Krylov theorem for
(9) $F(D^{2}u)-\partial_{t}u=0\text{ in }Q_{1}=B_{1}\times(-1,0].$
with
(10) $\begin{cases}F(M)=\min\{F^{\cap}(M),F^{\cup}(M)\},\ \forall M\in\mathcal{S}_{n}\\ F(0)=0,\ F^{\cap},F^{\cup}\text{ are uniformly elliptic}\\ F^{\cap}\text{ is concave in $M$},\ F^{\cup}\text{ is convex in $M$}.\end{cases}$
We recall that the Evans-Krylov theorem allows to pass from $C^{1,1}$ to $C^{2,\alpha}$ regularity in the stationary case, see e.g. [8, Chapter 6].
Theorem 4.1.
Let $u\in C(Q_{1})$ be a viscosity solution to (9) with $F$ satisfying (10). Then for some universal $\tilde{\alpha}\in(0,1)$ depending on $n,\lambda,\Lambda$ we have
$\|u\|_{C^{2+\tilde{\alpha},1+\tilde{\alpha}/2}(\overline{Q}_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})}.$
where $C$ is a constant depending on $\lambda,\Lambda,n,\tilde{\alpha}$.
Remark 4.2.
Differently from the elliptic case, the condition $F(0)=0$ cannot be in general dropped, as outlined in [24, Remark p. 257].
As a corollary, an argument by L. Caffarelli [7] provides the maximal $C^{\alpha}$ regularity for the non-homogeneous equation.
Before proving the result we recall some standard facts, see [25] and [18]:
Theorem 4.3 (Krylov-Safonov parabolic Hölder regularity).
Let $u\in C(Q_{1})$ be a viscosity solution of (4). Then, for some $\alpha>0$ depending on $n,\lambda,\Lambda$ we have that $u\in C^{\alpha,\alpha/2}(Q_{\frac{1}{2}})$ and the following regularity estimate holds
$\|u\|_{C^{\alpha,\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})},$
where $C$ is a constant depending on $\lambda,\Lambda,n$.
This Hölder property is the best one can get for equations with bounded and measurable coefficients in the parabolic dimension $n+1\geq 2$ by Theorem 1.1 (see also the discussion below) of [33]. The previous estimates are the cornerstone to deduce higher regularity of first-order: the next result holds for viscosity solutions of fully nonlinear parabolic homogeneous equations without any other assumption on $F$ other than the uniform ellipticity, cf. Theorem 2.3 in [33] or Lemma 3 p. 257 in [24].
Theorem 4.4 (First-order space-time Hölder regularity).
Let $u\in C(Q_{1})$ be a viscosity solution of (4). Then, for some $\alpha>0$ depending on $n,\lambda,\Lambda$ we have that $\partial_{t}u,Du\in C^{\alpha,\alpha/2}(Q_{\frac{1}{2}})$ and
$\|\partial_{t}u\|_{C^{\alpha,\alpha/2}(Q_{\frac{1}{2}})}+\|Du\|_{C^{\alpha,\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})},$
where $C$ depends on $\lambda,\Lambda,n$.
We also have the following interior a priori $C^{2,\beta}$ estimate for the stationary non-homogeneous equation driven by operators satisfying (10), see Corollary 1.3 and Remark 1.4 of [5].
Theorem 4.5.
Let $u\in C(B_{1})$ be a viscosity solution to $F(D^{2}u)=f(x)$ in $B_{1}$ with $f\in C^{\beta}(B_{1})$ for some $0<\beta<\bar{\beta}$, $\bar{\beta}\in(0,1)$ being a universal constant, and $F$ satisfying (10). Then, the following estimate holds
$\|u\|_{C^{2,\beta}(\overline{B}_{\frac{1}{2}})}\leq C(\|u\|_{L^{\infty}(B_{1})}+\|f\|_{C^{\beta}(B_{\frac{3}{4}})}),$
where $C$ depends on $n,\lambda,\Lambda,\beta$.
Next we exploit the $C^{2,\beta}_{x}$ estimate and the evolution equation to improve the time-Hölder exponent of first-order derivatives from Theorem 4.4 via a comparison principle argument.
Proposition 4.6.
Suppose that there exists $\beta\in(0,1)$ and a constant $C>0$ such that
$|D^{2}u(x,\cdot)-D^{2}u(y,\cdot)|\leq C|x-y|^{\beta},\ x,y\in B_{\frac{1}{2}}.$
Any $C^{2,1}$ solution to the uniformly parabolic problem (9) (without any other structural assumption on $F$) satisfies the estimate
$|Du(\cdot,t)-Du(\cdot,s)|\leq C^{\prime}|t-s|^{\frac{1+\beta}{2}}\text{ in }B_{\frac{1}{16}}.$
where $C^{\prime}$ is a positive constant depending on $C,n,\bar{\beta},\lambda,\Lambda$.
Proof.
The result is formally a consequence of Theorem 2.1 with $K,f=0$, after observing that for $e\in\mathbb{R}^{n}$, $|e|=1$,
$u_{h}(x,t):=\frac{u(x+he,t)-u(x,t)}{h}\in\mathcal{S}(\lambda/n,\Lambda),$
i.e. $u_{h}$ is a solution to a nondivergence form equation in a smaller cylinder. The assumptions on the regularity of $D^{2}u$ in space (for fixed times) imply that $u_{h}\in C^{1,\beta}$ uniformly with respect to $h$ and the result follows.
We give an explicit direct proof, again using the maximum principle at a higher order level, and the equivalence of the Hölder seminorms with those given by Taylor polynomials [23, Sections 3.3 and 8.5]. The latter suggests the form of the barrier function needed to run the foregoing argument. We have by Theorem 3.3.1 in [23]
$|u_{h}(z^{\prime},t)-u_{h}(z,t)-Du_{h}(z,t)\cdot(z^{\prime}-z)|\leq C|z^{\prime}-z|^{1+\beta}\text{ on }Q_{\frac{1}{2}-h}.$
We apply the Young’s inequality to find
$C|z^{\prime}-z|^{1+\alpha}\leq\varepsilon+\widetilde{C}\varepsilon^{-\frac{1-\beta}{1+\beta}}|z^{\prime}-z|^{2}.$
This implies the inequality
$u_{h}(z^{\prime},t)\leq u_{h}(z,t)+Du_{h}(z,t)\cdot(z^{\prime}-z)+\varepsilon+\widetilde{C}\varepsilon^{-\frac{1-\beta}{1+\beta}}|z^{\prime}-z|^{2}.$
We define
$\Phi^{+}(z^{\prime},t^{\prime})=u_{h}(z,t)+Du_{h}(z,t)\cdot(z^{\prime}-z)+\varepsilon+\widetilde{C}\varepsilon^{-\frac{1-\beta}{1+\beta}}|z^{\prime}-z|^{2}+\nu(t^{\prime}-t).$
It is immediate to check that, after choosing $\nu=\widetilde{C}n\Lambda\varepsilon^{-\frac{1-\beta}{1+\beta}}$, we have
$\mathcal{M}^{+}_{\lambda,\Lambda}(D^{2}\Phi^{+})-\partial_{t}\Phi^{+}\leq 0,$
so $\Phi^{+}$ is a classical supersolution to the equation solved by $u_{h}$ in $Q_{\frac{1}{4}}$ provided that $h<\frac{1}{8}$. Recall that $u_{h}\in\mathcal{S}(\lambda/n,\Lambda)$ and thus solves a linear parabolic homogeneous equation in nondivergence form. Choosing $\varepsilon$ small we have that $\Phi^{+}\geq u_{h}$ on $\partial B_{\frac{1}{4}}\times(-\frac{1}{16},0]$. By the maximum principle $\Phi^{+}\geq u_{h}$ in the interior of the cylinder. Taking $z^{\prime}=z$ and optimizing with respect to $\varepsilon$ we conclude
$u_{h}(z,t^{\prime})\leq u_{h}(z,t)+\bar{C}(t^{\prime}-t)^{\frac{1+\beta}{2}},$
for a positive constant $\bar{C}$. We can get a similar estimate from below by constructing a subsolution, say $\Phi^{-}$, to the PDE
$\mathcal{M}^{-}_{\lambda,\Lambda}(D^{2}\Phi^{-})-\partial_{t}\Phi^{-}\geq 0$
and applying the minimum principle. ∎
We are now ready with the proof of the main result:
Proof of Theorem 4.1.
Theorem 4.4 shows that viscosity solutions of (9) satisfy $\partial_{t}u\in C^{\tilde{\alpha},\tilde{\alpha}/2}$ for some small $\tilde{\alpha}\in(0,1)$ depending on $\lambda,\Lambda$ and $n$. Recall that such a result does not depend on the structure of the operator. Therefore, we can consider
$F(D^{2}u)=\partial_{t}u\in C^{\tilde{\alpha},\tilde{\alpha}/2}.$
We freeze the time variable and apply Theorem 4.5 on each time slice to find
$|D^{2}u(x,\cdot)-D^{2}u(y,\cdot)|\leq C|x-y|^{\tilde{\alpha}}.$
Applying Proposition 4.6 with $\beta=\tilde{\alpha}$ and the space regularity of $D^{2}u$ we conclude that
$|Du(\cdot,t)-Du(\cdot,s)|\leq C|t-s|^{\frac{1+\tilde{\alpha}}{2}}.$
Therefore, $D^{2}u$ is time-Hölder continuous with exponent $\tilde{\alpha}/2$ after invoking the interpolation argument from Lemma 3.1 p. 80 of [26] applied to $Du$. We recall briefly that it says that if a function $v$ is $C^{\gamma_{1}}_{t}$ and $Dv\in C_{x}^{\gamma_{2}}$, $\gamma_{1},\gamma_{2}\in(0,1)$, then $Dv\in C^{\gamma_{1}\gamma_{2}/(1+\gamma_{2})}_{t}$. This implies that $D^{2}u\in C^{\tilde{\alpha},\tilde{\alpha}/2}$.
Remark 4.7.
It is worth comparing the strategy outlined above with those already appeared in the literature of fully nonlinear parabolic equations. There are essentially two ways to achieve $C^{2+\alpha,1+\alpha/2}$ estimates for homogeneous fully nonlinear parabolic equations. The first one is to prove Evans-Krylov estimates directly for parabolic equations, following e.g. [22]. The second one is to consider (9) as $F(D^{2}u)=\partial_{t}u$, give an estimate of $\partial_{t}u$ and then consider the PDE for any fixed $t$ as an elliptic one: this viewpoint however requires the development of a $C^{2,\alpha}$ theory for the non-homogeneous elliptic equation $F(D^{2}u)=f(x)\in C^{\alpha}$. The approach carried out in Chapter 12 of [24] follows this second idea, and exploits an interpolation argument between $D^{2}u(\cdot,t)\in C^{\alpha}_{x}$ and $\partial_{t}u(\cdot,t)\in C^{\alpha}_{x}$ for fixed times $t$ (cf. Lemma 4 p. 259 in [24]), a different way with respect to our method. Related ideas were also employed in [1, 33] for parabolic equations in 3D.
It is worth pointing out that the foregoing approach applies to deduce interior $C^{2+\alpha,1+\alpha/2}$ bounds in the case of concave operators, with a different proof than Theorem 4.13 in [36] (in particular without using Lemma 4.16 therein). Here we provided a general strategy to deduce parabolic $C^{2+\alpha,1+\alpha/2}$ regularity through the maximum principle and the $C^{2,\alpha}$ regularity in the space variable only. We can generalize the previous argument providing an abstract result under the following hypothesis
  • (H)
    The stationary equation $F(D^{2}u)=f$, $f\in C^{\alpha}$, has interior $C^{2,\alpha}$ estimates for some $\alpha\in(0,1)$ with constant $c_{e}$.
Theorem 4.8.
Let $u\in C(Q_{1})$ be a viscosity solution of the equation $F(D^{2}u)-\partial_{t}u=0$ in $Q_{1}$. Assume that $F$ is uniformly elliptic and that (H) holds. Then, $u\in C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})$ and the following estimate holds
$\|u\|_{C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})}\leq C$
for a constant $C$ depending on $n,\lambda,\Lambda,\alpha,c_{e}$.
We conclude by discussing the case of non-homogeneous equations. Schauder’s estimates proved by L. Caffarelli [8, Section 8.1], see also [36], would imply $C^{2+\alpha,1+\alpha}$ estimates for the following equation
$F(x,t,D^{2}u)-\partial_{t}u=f(x,t)$
under the following assumptions on $F$:
  • (i)
    For every $(x_{0},t_{0})\in Q_{1}$, the operator $F(x_{0},t_{0},\cdot)$ is the minimum of a concave and a convex operator (possibly depending on $(x_{0},t_{0})$);
  • (ii)
    $F(\cdot,M)$ and $f(\cdot)$ are Hölder continuous functions with respect to the parabolic distance.
This implies that $F(x,t,D^{2}u)$ can be an Isaacs operator of the form (3) with variable Hölder continuous coefficients with respect to the parabolic distance.