### v3

# Interpolated time-Hölder regularity of solutions of fully nonlinear parabolic equations

Abstract.
We show interior Schauder estimates for a special class of fully nonlinear parabolic Isaacs equations by the maximum principle, providing an Evans-Krylov result for the model equation $\min\{\inf_{\beta}L_{\beta}u,\sup_{\gamma}L_{\gamma}u\}-\partial_{t}u=0$, where $L_{\beta},L_{\gamma}$ are linear operators with possibly variable Hölder coefficients. We also give a proof of the Evans-Krylov theorem for fully nonlinear uniformly parabolic equations for which a regularity theory of the stationary non-homogeneous equation is available.
Key words and phrases:
Evans-Krylov theorem, Fully nonlinear parabolic equations, Isaacs equations
2020 Mathematics Subject Classification:
35B65
The author is member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). He was partially supported by the INdAM-GNAMPA projects 2024 and 2025, the King Abdullah University of Science and Technology (KAUST) project CRG2021-4674 “Mean-Field Games: models, theory and computational aspects” and by the project funded by the EuropeanUnion – NextGenerationEU under the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.1 - Call PRIN 2022 No. 104 of February 2, 2022 of Italian Ministry of University and Research; Project 2022W58BJ5 (subject area: PE - Physical Sciences and Engineering) “PDEs and optimal control methods in mean field games, population dynamics and multi-agent models”
1. Introduction
In this note we provide a maximum principle approach to study interior maximal $C^{\alpha}$ smoothness (i.e. local $C^{2,\alpha}$ regularity with respect to the parabolic distance) of solutions to the following fully nonlinear parabolic model
(1) $F(D^{2}u)-\partial_{t}u=0\text{ in }Q_{1}:=B_{1}\times(-1,0]\subseteq\mathbb{R}^{n+1},$
where we assume
[…]
The uniform ellipticity for an operator $G=G(M):\mathrm{Sym}_{n}\to\mathbb{R}$, $\mathrm{Sym}_{n}$ being the space of $n\times n$ symmetric matrices, reads as
$\lambda\|N\|\leq G(M+N)-G(M)\leq\Lambda\|N\|,\ \forall M\in\mathrm{Sym}_{n},\ N\geq 0,\ \lambda\leq\Lambda,\ \|N\|=\sup_{|x|=1}|Nx|.$
Note that the uniform ellipticity of $F^{\cap},F^{\cup}$ implies that $F$ itself is uniformly elliptic. In this case we will call (1) uniformly parabolic. For instance, the result applies to the evolution PDE
(3) $\min\left\{\inf_{a\in\mathcal{A}}L_{a}u,\sup_{b\in\mathcal{B}}L_{b}u\right\}-\partial_{t}u=0,$
where $\mathcal{A},\mathcal{B}$ are arbitrary sets and $L_{a},L_{b}$ are linear operators of the form $L_{k}u=a_{ij}^{k}\partial_{ij}u+c_{k}$, $c_{k}=L_{k}0\in\mathbb{R}$, $a_{ij}^{k}$ with constant coefficients. This class can be written in the Isaacs form $\inf_{\sigma\in\Sigma}\sup_{\theta\in\Theta}L_{\sigma\theta}u-\partial_{t}u=0$, where the convex operators $\left(\sup_{\theta\in\Theta}L_{\sigma\theta}u\right)_{\sigma\in\Sigma}$ are all linear except for at most one $\sigma$. A prototype equation is the flow driven by the 3-operator
$F_{3}(D^{2}u)-\partial_{t}u=\min\{L_{1}u,\max\{L_{2}u,L_{3}u\}\}-\partial_{t}u,$
where $L_{1},L_{2},L_{3}$ are linear uniformly elliptic operators with constant coefficients. Our result provides a time-dependent counterpart of the maximal $C^{\alpha}$ regularity found by X. Cabré and L. Caffarelli [5] and it can be regarded as an Evans-Krylov type result for a (time-dependent) nonconvex/nonconcave equation. We recall that the Evans-Krylov property means the boost of regularity
$C^{2,1}\rightarrow C^{2+\alpha,1+\alpha/2},$
see Theorem 6.1 of [6] for the elliptic result.
The outline of the proof of the maximal parabolic $C^{\alpha}$ regularity under the assumption (2) is the following: one first proves that for the homogeneous equation $F(D^{2}u)-\partial_{t}u=0$ we have
$\partial_{t}u,Du\in C^{\alpha,\alpha/2}$
for some small universal $\alpha\in(0,1)$, by the Krylov-Safonov theory [20] (here $C_{t}^{\alpha/2}$ denotes the Hölder continuity in time, uniformly in space). Indeed, the first derivative (or its incremental quotient), both in time and space, satisfies a linear nondivergence form equation, since the equation is invariant in space-time. This step does not require neither convexity nor concavity assumptions on $F$, and it holds for continuous viscosity solutions of any uniformly parabolic equation, see Theorem 3.4. Then, for each time slice, which we freeze and do not display, we write
$F(D^{2}u)=g$
where $g=\partial_{t}u\in C^{\alpha,\alpha/2}$. This implies by the stationary (space) regularity of [5]
$u\in C^{2,\alpha}_{x}$
at each time layer. To reach parabolic $C^{2+\alpha,1+\alpha/2}$ regularity of the solution to the parabolic equation (and in particular $\alpha/2$ time-regularity of second derivatives) we exploit a maximum principle method. More precisely, exploiting the information $D^{2}u\in C^{\alpha}_{x}$, we improve the time-regularity of the gradient showing that
$Du\in C_{t}^{\frac{1+\alpha}{2}}$
via the comparison principle. We then exploit an interpolation argument between $Du\in C_{t}^{(1+\alpha)/2}$ and $D^{2}u\in C^{\alpha}_{x}$ to conclude that $D^{2}u\in C^{\alpha/2}_{t}$, proving thus the claimed $C^{2+\alpha,1+\alpha/2}$ regularity. Note that a direct interpolation with the sole information $\partial_{t}u,Du\in C^{\alpha,\alpha/2}$ available from the Krylov-Safonov theory fails to yield the expected $\alpha/2$-Hölder regularity in time of $D^{2}u$. However, it is enough in general to reach classical smoothness for a smaller exponent $\bar{\alpha}=\frac{\alpha}{2+\alpha}<\frac{\alpha}{2}$.
In the case of the heat equation, a maximum principle approach was used to prove the time-Hölder regularity of the Hessian in [15] via an argument of A. Brandt [2].
Notably, this method provides a new proof of the parabolic $C^{2+\alpha,1+\alpha/2}$ regularity for fully nonlinear uniformly parabolic equations knowing the stationary $C^{2,\alpha}$ estimate for nonhomogeneous equations via the maximum principle; see the end of Section 3. This is the case of fully nonlinear uniformly parabolic equations in two space dimensions [1, 28], operators with Cordes conditions [13, 10], concave [11, 19] and quasiconcave equations [13], operators of twisted type [7]. A comparison with other approaches to study the parabolic regularity of fully nonlinear equations is in Remark 3.9.
The study of these estimates are motivated by the challenging open question of finding some special structures of $F$ (intermediate among the sole uniform ellipticity and weaker than convexity) guaranteeing classical smoothness of solutions when the parabolic dimension $n+1\geq 4$, cf. the introduction of [13] and Section 2 for references on the subject.
Following [5], we mention that these interior estimates could be used to prove an existence result for the Dirichlet problem by a refinement of the method of continuity based on weighted interior bounds. This approach was first introduced by J. H. Michael [25] for linear stationary PDEs and extended to linear parabolic equations in [30]. We do not pursue this direction here, but just recall that the application of the classical method of continuity requires estimates up to the boundary along with a smooth operator, namely $F$ must be at least of class $C^{1}$. Moreover, we emphasize that the process of smoothing fully nonlinear operators is known only for concave/convex equations, and hence one needs a different approach for other classes of non-smooth operators.
Plan of the paper. Section 2 provides a brief survey of the known literature of the $C^{2,\alpha}$ solvability of fully nonlinear parabolic equations. Section 3 is devoted to the application of the maximum principle approach to obtain $C^{2+\alpha,1+\alpha/2}$ regularity for a time-dependent counterpart of an Isaacs model introduced by X. Cabré and L. Caffarelli. We finally state an abstract result that provides $C^{2+\alpha,1+\alpha/2}$ regularity provided that $C^{2,\alpha}$ estimates for the elliptic problem are available.
Notation
We denote by $B_{r}(x)$ the ball of center $x$ and radius $r$. When $x_{0}=0$ we simply write $B_{r}$.
$Q_{r}(x,t)$ is the parabolic cylinder $B_{r}(x)\times(t-r^{2},t)$, and we write $Q_{r}$ when $(x,t)=(0,0)$.
We denote, given an open set $\Omega$, by $\partial_{\mathrm{par}}(\Omega\times(a,b))$ the parabolic boundary of a set.
We denote by $|u|_{0;\Omega}$ the sup-norm of $u$, i.e. $\|u\|_{C(\Omega)}$ (both in the elliptic and the parabolic case).
Let $Q\subset\Omega\times(0,T)$ and $\alpha\in(0,1)$. We denote by $d((x,t),(y,s))=|x-y|+|t-s|^{\frac{1}{2}}$ the standard parabolic distance and
  • $C^{\alpha,\alpha/2}(Q)$, $\alpha\in(0,1]$, the space of those $u:Q\to\mathbb{R}$ such that
    $\|u\|_{C^{\alpha,\alpha/2}(Q)}:=|u|_{0;Q}+[u]_{C^{\alpha,\alpha/2}(Q)}=|u|_{0;Q}+\sup_{\begin{subarray}{c}(x,t),(y,s)\in Q,\\ (x,t)\neq(y,s)\end{subarray}}\frac{|u(x,t)-u(y,s)|}{d^{\alpha}((x,t),(y,s))};$
  • $C^{1+\alpha,(1+\alpha)/2}(Q)$ the space of functions $u$ whose spatial gradient exists classically and equipped with the norm
    $\|u\|_{C^{1+\alpha,(1+\alpha)/2}(Q)}:=|u|_{0;Q}+|Du|_{0;Q}+\sup_{\begin{subarray}{c}(x,t),(y,s)\in Q,\\ (x,t)\neq(y,s)\end{subarray}}\frac{|u(x,t)-u(y,s)|}{d^{1+\alpha}((x,t),(y,s))}.$
    In particular, any $u\in C^{1+\alpha,(1+\alpha)/2}(Q)$ is such that each component of $Du$ belongs to $C^{\alpha,\alpha/2}(Q)$ and $u$ is Hölder continuous with exponent $(1+\alpha)/2$ in the time variable;
  • $C^{2+\alpha,1+\alpha/2}(Q)$ the space of functions $u$ such that
    $\|u\|_{C^{2+\alpha,1+\alpha/2}(Q)}:=|u|_{0;Q}+\sum_{i=1}^{n}\|\partial_{x_{i}}u\|_{C^{1+\alpha,(1+\alpha)/2}(Q)}+\|\partial_{t}u\|_{C^{\alpha,\alpha/2}(Q)}$
    This is equivalent to say that $D^{2}u$ belongs to $C^{\alpha,\alpha/2}(Q)$ and $\partial_{t}u$ belongs to $C^{\alpha,\alpha/2}(Q)$: it is a consequence of Remark 8.8.7 in [18]. In this case we can consider the space $C^{2+\alpha,1+\alpha/2}(Q)$ equipped with the norm
    $\|u\|_{C^{2+\alpha,1+\alpha/2}(Q)}:=|u|_{0;Q}+|Du|_{0;Q}+\|D^{2}u\|_{C^{\alpha,\alpha/2}(Q)}+\|\partial_{t}u\|_{C^{\alpha,\alpha/2}(Q)}.$
    For more properties on these spaces we refer to [18]. Furthermore, we denote by
    $\|u\|_{C^{2,1}(Q)}:=\sum_{2i+j\leq 2}|\partial_{t}^{i}D_{x}^{j}u|_{0;Q}.$
    We will also use the equivalence (see e.g. p. 120 of [18]) between the Hölder seminorm $u]_{C^{\alpha,\alpha/2}(Q_{1})}$ previously defined and the seminorm
    $[u]^{\prime}_{C^{\alpha,\beta}(Q_{1})}$ $=\sup_{t\in(-1,0)}\sup_{\begin{subarray}{c}x,y\in B_{1},\\ x\neq y\end{subarray}}\frac{|u(x,t)-u(y,t)|}{|x-y|^{\alpha}}+\sup_{x\in B_{1}}\sup_{\begin{subarray}{c}t,s\in(-1,0),\\ t\neq s\end{subarray}}\frac{|u(x,t)-u(x,s)|}{|t-s|^{\beta}}$
    $=\sup_{t\in(-1,0)}[u(\cdot,t)]_{C^{\alpha}(B_{1})}+\sup_{x\in B_{1}}[u(x,\cdot)]_{C^{\beta}((-1,0))}.$
2. A survey on the (higher) regularity theory of fully nonlinear parabolic equations
In this section we survey on regularity results at the level of Hölder spaces for fully nonlinear parabolic equations, focusing on those of the form
(4) $F(D^{2}u)-\partial_{t}u=0\text{ in }B_{1}\times(-1,0]\subset\mathbb{R}^{n+1}.$
We will in particular concentrate on structural conditions, possibly involving the dimension, guaranteeing classical solutions. The interested reader can find a complete updated account in the book [12] for the elliptic theory. We do not aim at discussing here $W^{2,1}_{q}$ regularity of solutions and we refer to [9, 19, 31] for more details and to the next section for the main results on the low-regularity theory.
The simplest result for (4) says that solutions to uniformly parabolic equations in $D=2+1$ (here we denote by $D=n+1$ the dimension in space-time) are always classical without any other assumption on $F$, see [1, 13, 28] and the references therein. Its elliptic analogue was proved by L. Nirenberg (in the elliptic dimension $n=2$) [12, Theorem 4.9], and its proof, based on the De Giorgi-Nash-Moser theory applied to the equation solved by second derivatives, does not seem to generalize to parabolic equations. It is also important to mention a result of S. Kruzhkov showing that the same result holds in the $D=2$ case ($n=1$ variable in space), cf. [16] and Section XIV.7 in [22].
Theorem 2.1.
Assume that $F$ is uniformly elliptic and $n=2$. Then viscosity solutions to (4) (in dimension $n+1=3$) are always classical and belong to $C^{2+\alpha,1+\alpha/2}_{\mathrm{loc}}$ for some small universal $\alpha\in(0,1)$. It holds
$\|u\|_{C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})}.$
A second result holds under Cordes-type assumptions on the ellipticity in any space dimension, see Theorem 6.4 in [13] or [10].
Theorem 2.2.
Assume that $F$ is uniformly elliptic and $\frac{\Lambda}{\lambda}\leq 1+\delta$ for a small (explicit) constant $\delta$ depending only on $n$. Then viscosity solutions to (4) are always classical (in any dimension) and belong to $C^{2+\alpha,1+\alpha/2}_{\mathrm{loc}}$ for some small universal $\alpha$. In addition, we have the regularity estimate
$\|u\|_{C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})}.$
We also mention that $C^{2+\alpha,1+\alpha/2}$ estimates can be obtained without any concavity assumption for short time horizons, see e.g. Chapter 8 of [24], [23] and the references therein. More recently, maximal $C^{\alpha}$ estimates were also achieved for flat solutions (i.e. with small $L^{\infty}$ norm) in [32] without concavity assumptions on $F$, extending a result due to O. Savin for stationary equations. These latter and the previous results are true for any uniformly parabolic equation, and hence for uniformly parabolic Isaacs equations given by
$\inf_{\eta\in A}\sup_{\gamma\in B}\{L_{\eta\gamma}u-f_{\eta\gamma}\}=0,$
under appropriate restrictions (e.g. on the dimension, the coefficients, the solution or the time horizon). We also emphasize that some counterexamples to the smoothness of solutions when $F$ is only uniformly parabolic can be found in the recent paper [28], which provides a time-dependent counterpart of the analysis in [26].
A counterpart of the result by Evans-Krylov [11, 19] provides parabolic $C^{2,\alpha}$ estimates under concavity assumptions on $F$, cf. [31], see also [22, 17, 19] for smooth operators.
Theorem 2.3.
Assume that $F$ is uniformly elliptic and concave. Then viscosity solutions to (4) are always classical and belong to $C^{2+\alpha,1+\alpha/2}_{\mathrm{loc}}$ for some small universal $\alpha$. It holds
$\|u\|_{C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})}.$
Concavity on the second derivatives was recently weakened to the requirement of convexity of the superlevel sets, cf. [13]. The latter also contains some $C^{1,1}$ estimates under concavity-type conditions at infinity implying, among others, $W^{2,1}_{q}$ estimates for non-homogeneous equations. A recent account on the treatment of fully nonlinear parabolic equations with relaxed concavity conditions can be found in [19].
We also mention that slightly before the Evans-Krylov theory, the paper [27] provided an obstacle problem approach to reach parabolic $C^{2,\beta}$, $\beta\in(0,1)$ small and universal, regularity for the model two-operator equation
$\max\{\partial_{t}u-L_{1}u,\partial_{t}u-L_{2}u\}=0$
(here $L_{1},L_{2}$ are uniformly elliptic linear operators) by reducing the problem to a variational inequality, an idea introduced by Brézis-Evans [4] for stationary equations. More references on the early regularity theory of fully nonlinear elliptic and parabolic equations can be found in [22, p. 382-384]. We are not aware of any nonconcave/nonconvex $F$ guaranteeing $C^{2+\alpha,1+\alpha/2}$ estimates in the time-dependent framework, except for [29] and the aforementioned papers dealing with certain restrictions on the data of the problem.
3. $C^{2+\alpha,1+\alpha/2}$ regularity for fully nonlinear parabolic Isaacs equations
Our main result is an Evans-Krylov theorem for
(5) $F(D^{2}u)-\partial_{t}u=0\text{ in }Q_{1}=B_{1}\times(-1,0].$
with
[…]
We recall that the Evans-Krylov theorem allows us to pass from $C^{1,1}$ to $C^{2,\alpha}$ regularity in the stationary case, see e.g. [6, Chapter 6].
Theorem 3.1.
Let $u\in C(Q_{1})$ be a viscosity solution to (5) with $F$ satisfying (6). Then for some universal $\tilde{\alpha}\in(0,1)$ depending on $n,\lambda,\Lambda$ we have
$\|u\|_{C^{2+\tilde{\alpha},1+\tilde{\alpha}/2}(\overline{Q}_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})}.$
where $C$ is a constant depending on $\lambda,\Lambda,n,\tilde{\alpha}$.
Remark 3.2.
Differently from the elliptic case, the condition $F(0)=0$ cannot be in general dropped, as outlined in [19, Remark p. 257].
As a corollary, an argument by L. Caffarelli [6, Section 8.1], see also [31, Theorem 1.1] for the time-dependent case, provides the maximal $C^{\alpha}$ regularity, $\alpha<\tilde{\alpha}$, for the non-homogeneous equation
$F(x,t,D^{2}u)-\partial_{t}u=f(x,t)$
under the following assumptions on $F$:
  • (i)
    For every $(x_{0},t_{0})\in Q_{1}$, the operator $F(x_{0},t_{0},\cdot)$ is the minimum of a concave and a convex operator (possibly depending on $(x_{0},t_{0})$);
  • (ii)
    $F(\cdot,M)$ and $f(\cdot)$ are Hölder continuous functions with exponent $\alpha$ with respect to the parabolic distance.
This implies that $F(x,t,D^{2}u)$ can be an Isaacs operator of the form (3) with variable Hölder continuous coefficients with respect to the parabolic distance.
We denote by $\mathcal{M}^{\pm}_{\lambda,\Lambda}$ the Pucci’s extremal operators with constants $0<\lambda\leq\Lambda$ defined by
$\mathcal{M}^{+}_{\lambda,\Lambda}(M)=\sup_{\lambda I_{n}\leq A\leq\Lambda I_{n}}\mathrm{Tr}(AM)=\Lambda\sum_{e_{k}>0}e_{k}+\lambda\sum_{e_{k}<0}e_{k},$
$\mathcal{M}^{-}_{\lambda,\Lambda}(M)=\sup_{\lambda I_{n}\leq A\leq\Lambda I_{n}}\mathrm{Tr}(AM)=\lambda\sum_{e_{k}>0}e_{k}+\Lambda\sum_{e_{k}<0}e_{k},$
where $e_{k}=e_{k}(M)$ are the eigenvalues of $M$. Moreover, $\underline{\mathcal{S}}(\lambda,\Lambda,f)$ stands for the space of continuous functions $u$ in $\Omega$ that are subsolutions of the evolutive Pucci’s maximal equation
$\mathcal{M}^{+}_{\lambda,\Lambda}(D^{2}u)-\partial_{t}u=f(x,t)\text{ in the viscosity sense in $\Omega$},$
Similarly, $\overline{\mathcal{S}}(\lambda,\Lambda,f)$ will denote the space of continuous functions $u$ in $\Omega$ that are supersolutions to the Pucci’s minimal equation
$\mathcal{M}^{-}_{\lambda,\Lambda}(D^{2}u)-\partial_{t}u=f(x,t)\text{ in the viscosity sense in $\Omega$}.$
Solutions of fully nonlinear uniformly elliptic equations $F(D^{2}u)-\partial_{t}u=0$ belong to the class $\mathcal{S}(\lambda,\Lambda,0)=\overline{\mathcal{S}}(\lambda,\Lambda,0)\cap\underline{\mathcal{S}}(\lambda,\Lambda,0)$.
Before proving Theorem 3.1 we recall some standard facts, see [20] and [14]:
Theorem 3.3 (Krylov-Safonov parabolic Hölder regularity).
Let $u\in C(Q_{1})$ be a solution of the viscosity inequalities
[…]
Then, for some $\alpha>0$ depending on $n,\lambda,\Lambda$ we have that $u\in C^{\alpha,\alpha/2}(Q_{\frac{1}{2}})$ and the following regularity estimate holds
$\|u\|_{C^{\alpha,\alpha/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})},$
where $C$ is a constant depending on $\lambda,\Lambda,n$.
This Hölder property is the best one can get for equations with bounded and measurable coefficients in the parabolic dimension $D=n+1\geq 2$ by [28, Theorem 2.2 and the discussion below]. The previous estimates are the cornerstone to deduce higher regularity of first-order: the next result holds for viscosity solutions of any fully nonlinear parabolic homogeneous equations without any other assumption on $F$ other than the uniform ellipticity, cf. Theorem 2.3 in [28], Theorems 4.8 and 4.9 in [31] or Lemma 3 p. 257 in [19] (see also Corollary 5.7 in [6] and [12, Theorem 4.24] for the elliptic case).
Theorem 3.4 (First-order space-time Hölder regularity).
Let $u\in C(Q_{1})$ be a viscosity solution of $(D^{2}u)-\partial_{t}u=0, FF$ uniformly elliptic. Then, for some $\theta>0$ depending on $n,\lambda,\Lambda$ we have that $\partial_{t}u,Du\in C^{\theta,\theta/2}(Q_{\frac{1}{2}})$ and
$\|\partial_{t}u\|_{C^{\theta,\theta/2}(Q_{\frac{1}{2}})}+\|Du\|_{C^{\theta,\theta/2}(Q_{\frac{1}{2}})}\leq C\|u\|_{L^{\infty}(Q_{1})},$
where $C$ depends on $\lambda,\Lambda,n$.
We now state the main result we need to prove parabolic maximal $C^{\alpha}$ regularity for (5). This is an interior $C^{2,\beta}$ estimate, $\beta$ small and universal, for the stationary non-homogeneous equation driven by operators satisfying (6), see Corollary 1.3 and Remark 1.4 of [5].
Theorem 3.5.
Let $u\in C(B_{1})$ be a viscosity solution to $F(D^{2}u)=f(x)$ in $B_{1}$ with $f\in C^{\gamma}(B_{1})$ for some $0<\gamma<\bar{\gamma}$, $\bar{\gamma}\in(0,1)$ being a universal constant, and $F$ satisfying (6). Then, the following estimate holds
$\|u\|_{C^{2,\gamma}(\overline{B}_{\frac{1}{2}})}\leq C(\|u\|_{L^{\infty}(B_{1})}+\|f\|_{C^{\gamma}(\overline{B}_{\frac{3}{4}})}),$
where $C$ depends on $n,\lambda,\Lambda,\gamma$.
Remark 3.6.
The universal exponent $\bar{\gamma}\in(0,1)$ is the Hölder regularity exponent of second derivatives of solutions solving the constant coefficient equation, where $F$ is uniformly elliptic and satisfies (6).
Next we exploit the above $C^{2,\gamma}_{x}$ estimate and the evolution equation to improve the time-Hölder exponent of first-order derivatives from Theorem 3.4 via a comparison principle argument. This is inspired by [1].
Proposition 3.7.
Suppose that there exist $\beta\in(0,1)$ and a constant $C>0$ such that
$|D^{2}u(x,\cdot)-D^{2}u(y,\cdot)|\leq C|x-y|^{\beta},\ x,y\in B_{\frac{1}{2}}.$
Any $C^{2,1}$ solution to the uniformly parabolic equation (5) (without any other structural assumption on $F$) satisfies the estimate
[…]
where $C^{\prime}$ is a positive constant depending on $C,n,\beta,\lambda,\Lambda$.
Remark 3.8.
Note that the validity of the Hölder regularity condition on the second derivatives hides some additional structural conditions on $F$ beyond the uniform ellipticity, unless the space dimension $n=1,2$.
Proof.
First note that for $e\in\mathbb{R}^{n}$, $|e|=1$,
$u_{h}(x,t):=\frac{u(x+he,t)-u(x,t)}{h}\in\mathcal{S}(\lambda/n,\Lambda),$
i.e. $u_{h}$ is a solution to a nondivergence form equation in a smaller cylinder. This means (see e.g. [6, Proposition 5.5] or [31, Theorem 4.6]) that we have the validity of the following inequalities
(7) $\mathcal{M}^{+}_{\lambda,\Lambda}(D^{2}u_{h})-\partial_{t}u_{h}\geq 0\text{ in }Q_{\frac{1}{2}-h}$
and
(8) $\mathcal{M}^{-}_{\lambda,\Lambda}(D^{2}u_{h})-\partial_{t}u_{h}\leq 0\text{ in }Q_{\frac{1}{2}-h}.$
Moreover, the assumptions on the regularity of $D^{2}u$ in space (for fixed times) imply that $u_{h}\in C^{1,\beta}_{x}$, $\beta\in(0,1)$, uniformly with respect to $h$ .
We give a proof using the maximum principle and the equivalence of the Hölder seminorms with those given by Taylor polynomials [18, Sections 3.3 and 8.5]. We have by Theorem 3.3.1 in [18]
$|u_{h}(z^{\prime},t)-u_{h}(z,t)-Du_{h}(z,t)\cdot(z^{\prime}-z)|\leq C|z^{\prime}-z|^{1+\beta}\text{ on }Q_{\frac{1}{2}-h}.$
We apply the weighted Young’s inequality to find for a positive constant $\widetilde{C}$
$C|z^{\prime}-z|^{1+\beta}\leq\varepsilon+\widetilde{C}\varepsilon^{-\frac{1-\beta}{1+\beta}}|z^{\prime}-z|^{2}.$
This implies the inequality
$u_{h}(z^{\prime},t)\leq u_{h}(z,t)+Du_{h}(z,t)\cdot(z^{\prime}-z)+\varepsilon+\widetilde{C}\varepsilon^{-\frac{1-\beta}{1+\beta}}|z^{\prime}-z|^{2}.$
We define
$\Phi^{+}(z^{\prime},t^{\prime})=u_{h}(z,t)+Du_{h}(z,t)\cdot(z^{\prime}-z)+\varepsilon+\widetilde{C}\varepsilon^{-\frac{1-\beta}{1+\beta}}|z^{\prime}-z|^{2}+\nu(t^{\prime}-t).$
One can check, after choosing $\nu\geq 2\widetilde{C}n\Lambda\varepsilon^{-\frac{1-\beta}{1+\beta}}$, the validity of
$\mathcal{M}^{+}_{\lambda,\Lambda}(D^{2}\Phi^{+})-\partial_{t}\Phi^{+}\leq 0,$
so $\Phi^{+}$ is a classical supersolution to the equation solved by $u_{h}$ in $Q_{\frac{1}{4}}$, provided that $h<\frac{1}{8}$. Recall that $u_{h}\in\mathcal{S}(\lambda/n,\Lambda)$ and thus solves a linear parabolic homogeneous equation in nondivergence form, see (7)-(8). Choosing $\varepsilon$ small enough, $\varepsilon<K\|u\|_{C^{2}}^{-\frac{1-\beta}{1+\beta}}$ for a suitable $K>0$, we have that $\Phi^{+}\geq u_{h}$ on $\partial B_{\frac{1}{4}}\times(-\frac{1}{16},0]$. By the maximum principle [8, Theorem 8.2] $\Phi^{+}\geq u_{h}$ in the interior of the cylinder. Taking $z^{\prime}=z$ and optimizing with respect to $\varepsilon$ we conclude
$u_{h}(z,t^{\prime})\leq u_{h}(z,t)+\bar{C}_{1}(t^{\prime}-t)^{\frac{1+\beta}{2}},$
for a positive constant $\bar{C}_{1}$. We can get a similar estimate from below by constructing a subsolution, say $\Phi^{-}$, to the PDE
$\mathcal{M}^{-}_{\lambda,\Lambda}(D^{2}\Phi^{-})-\partial_{t}\Phi^{-}\geq 0$
and applying the comparison principle. More precisely, we have
$u_{h}(z^{\prime},t)\geq u_{h}(z,t)+Du_{h}(z,t)\cdot(z^{\prime}-z)-\varepsilon-\widetilde{C}\varepsilon^{-\frac{1-\beta}{1+\beta}}|z^{\prime}-z|^{2}.$
We define
$\Phi^{-}(z^{\prime},t^{\prime})=u_{h}(z,t)+Du_{h}(z,t)\cdot(z^{\prime}-z)-\varepsilon-\widetilde{C}\varepsilon^{-\frac{1-\beta}{1+\beta}}|z^{\prime}-z|^{2}-\nu(t^{\prime}-t)$
and choose $\nu\geq 2\widetilde{C}n\lambda\varepsilon^{-\frac{1-\beta}{1+\beta}}$ to conclude that $\mathcal{M}^{-}_{\lambda,\Lambda}(D^{2}\Phi^{-})-\partial_{t}\Phi^{-}\geq 0$. By the maximum principle we have $\Phi^{-}\leq u_{h}$ in the interior of the cylinder, and proceeding as above we have for a positive constant $\bar{C}_{2}$ the estimate
$u_{h}(z,t^{\prime})\geq u_{h}(z,t)-\bar{C}_{2}(t^{\prime}-t)^{\frac{1+\beta}{2}}.$
We are now ready with the proof of the main result:
Proof of Theorem 3.1.
Theorem 3.4 shows that viscosity solutions of (5) satisfy $\partial_{t}u\in C^{\theta,\theta/2}$ for some small $\theta\in(0,1)$ depending on $\lambda,\Lambda$ and $n$ ($\theta$ is the exponent of Theorem 3.4). Recall that such a result does not depend on the structure of the operator and it requires only its uniform ellipticity. Therefore, we can freeze the time-variable and consider the solution $u=u(\cdot,t)$ of
$F(D^{2}u(\cdot,t))=\partial_{t}u(\cdot,t)\in C^{\theta}.$
Note that $\partial_{t}u(\cdot,t)$ is Hölder continuous with a possibly smaller exponent by the inclusion of Hölder spaces, namely $\partial_{t}u(\cdot,t)\in C^{\tilde{\alpha}}$, $\tilde{\alpha}\in(0,\bar{\gamma})\cap(0,\theta]$. By Theorem 3.5 we have that $D^{2}u(\cdot,t)$ exists and it is Hölder continuous with an exponent $\tilde{\alpha}\in(0,\bar{\gamma})\cap(0,\theta]$ on each time slice, namely
$|D^{2}u(x,\cdot)-D^{2}u(y,\cdot)|\leq C|x-y|^{\tilde{\alpha}}$.
Applying Proposition 3.7 with $\beta=\tilde{\alpha}$ we conclude that
$|Du(\cdot,t)-Du(\cdot,s)|\leq C|t-s|^{\frac{1+\tilde{\alpha}}{2}}.$
Therefore, $D^{2}u$ is time-Hölder continuous with exponent $\tilde{\alpha}/2$ after invoking the interpolation argument from Lemma 3.1 p. 80 of [21] applied to $Du$, see also [1, Section 3.4]. We recall briefly that it says that if a function $v$ is $C^{\gamma_{1}}_{t}$ and $Dv\in C_{x}^{\gamma_{2}}$, $\gamma_{1},\gamma_{2}\in(0,1)$, then $Dv\in C^{\gamma_{1}\gamma_{2}/(1+\gamma_{2})}_{t}$. This implies that $D^{2}u\in C^{\tilde{\alpha},\tilde{\alpha}/2}$.
Remark 3.9.
It is worth comparing the strategy outlined above with those already appeared in the literature of fully nonlinear parabolic equations. There are essentially two ways to achieve $C^{2+\alpha,1+\alpha/2}$ estimates for homogeneous fully nonlinear parabolic equations. The first one is to prove Evans-Krylov estimates directly for parabolic equations, as it is done in [17] for concave equations. The second one is to consider (5) as $F(D^{2}u)=\partial_{t}u$, give an estimate of $\partial_{t}u$ and then consider the PDE for any fixed $t$ as an elliptic one: this viewpoint however requires the development of a $C^{2,\alpha}$ theory for the non-homogeneous elliptic equation $F(D^{2}u)=f(x)\in C^{\alpha}$. The approach carried out in Chapter 12 of [19], see also [3, Appendix A], follows this second idea, and exploits an interpolation argument between $D^{2}u(\cdot,t)\in C^{\alpha}_{x}$ and $\partial_{t}u(\cdot,t)\in C^{\alpha}_{x}$ for fixed times $t$ (cf. Lemma 4 p. 259 in [19]), a slightly different way with respect to our method. Related ideas were also employed in [1, 28] for parabolic equations in 3D.
The foregoing approach applies to deduce interior $C^{2+\alpha,1+\alpha/2}$ bounds in the case of concave operators with a different proof than Theorem 4.13 in [31] (in particular without using Lemma 4.16 therein), [19, Theorem 12.2.1] and [3, Appendix A]. Here we provided a general strategy to deduce parabolic $C^{2+\alpha,1+\alpha/2}$ regularity through the maximum principle and the $C^{2,\alpha}$ regularity in the space variable only. We can generalize the previous argument providing an abstract result under the following hypothesis:
  • (H)
    The stationary non-homogeneous equation $F(D^{2}u)=f$, $f\in C^{\alpha}$, has interior $C^{2,\alpha}$ estimates for some universal $\alpha\in(0,1)$ with constant $c_{e}$.
Theorem 3.10.
Let $u\in C(Q_{1})$ be a viscosity solution of the equation $F(D^{2}u)-\partial_{t}u=0$ in $Q_{1}$. Assume that $F$ is uniformly elliptic, $F(0)=0$, and that (H) holds (i.e. the nonhomogeneous equation admits local $C^{2,\alpha}$ estimates). Then, $u\in C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})$ and the following estimate holds
$\|u\|_{C^{2+\alpha,1+\alpha/2}(Q_{\frac{1}{2}})}\leq C$
for a constant $C$ depending on $n,\lambda,\Lambda,\alpha,c_{e}$.
Proof.
Since $F$ is uniformly elliptic, we have $\partial_{t}u,Du\in C^{\alpha,\alpha/2}$. Then, the assumption (H) yields $D^{2}u\in C^{\alpha}_{x}$, uniformly in time by considering the equation $F(D^{2}u(\cdot,t))=\partial_{t}u(\cdot.t)$. Therefore, we can proceed as in the proof of Theorem 3.1 to prove that $D^{2}u\in C_{t}^{\alpha/2}$. ∎